# Optimal and Fair Encouragement Policy Evaluation and Learning

### v4

@@ 4.1 Case study: Text message reminders for SNAP recertification
Each year, households receiving SNAP (food assistance) must complete a short recertification interview to verify continued eligibility. Missing this interview results in termination of benefits, even for households that remain eligible.
@@
We study a pilot program in San Francisco in which clients could opt in to receive a text-message reminder about the recertification deadline. The reminder encouraged participants to complete the required interview—our effective treatment—so that their benefits would continue. Since many who fail to recertify are in fact eligible, increasing interview attendance often improves benefit continuity and household welfare.
@@
Using data from Homonoff and Somerville (2021), we reinterpret the reminder as an encouragement $R$ affecting the take-up of treatment $T$ (interview completion), which in turn influences the outcome $Y$ (next-year SNAP benefits). We focus on heterogeneity in both the reminder’s compliance effect and the treatment effect of interview completion, with particular attention to racial disparities in their alignment.
Text reminders are a low-cost operational lever that can reduce administrative burdens, but limited budgets may constrain how widely they are deployed. Our analysis evaluates how to target such reminders to maximize overall benefit take-up while considering fairness in access and impact across groups.
Descriptives: heterogeneity vs nudgeability.
In Figure˜1, we compare the distributions within each race (binarized to white $A=0$ or nonwhite $A=1$) of the heterogeneous treatment effect $\mathbb{E}[Y(1)-Y(0)\mid X],$ the heterogeneous compliance effect $p_{1\mid 1}(X)-p_{1\mid 0}(X),$ and the heterogeneous encouragement effect, which is the product of the previous two: $(\mathbb{E}[Y(1)-Y(0)\mid X])(p_{1\mid 1}(X)-p_{1\mid 0}(X))$. In a setting with fixed costs of recommendation and treatment, the default approach of viewing the recommendation as the treatment, i.e., an optimal unconstrained encouragement regime, sorts by the latter score.
In this setting, we find that treatment efficacy is broadly correlated with takeup: Higher treatment effects are rank correlated with higher compliance effects. However, the strength of this relationship, quantified via the Spearman’s $\rho_{SP}$ rank correlation coefficient, varies across groups, leading to inequality. Among white beneficiaries only, the rank correlation coefficient between heterogeneous treatment and compliance effects is 0.68, while it is 0.45 for nonwhite beneficiaries. While both correlations are statistically significantly positive, the relationship is weaker for nonwhite than for white beneficiaries, in particular with regard to those most likely to take up.
@@
A central concern here is that worse-off groups might have more to gain by accessing the treatment (i.e., have heterogeneous treatment effects of greater magnitude) but may respond less to nudges/recommendations into treatment (because of behavioral frictions, time/resource constraints, higher need, etc.). In this setting, decreasing marginal utilities of money establish the former, while the process of filling out the burdensome benefits application or making it to a recertification interview under time, childcare, transportation, or resource constraints speaks to the relative difficulty posed by the latter. The density plots in Figure˜1 underscore this phenomenon. The first plot shows the treatment effect heterogeneity: The distribution is wider for nonwhite than for white beneficiaries; i.e., it has more mass at higher magnitudes of $\tau$ and less at lower magnitudes of $\tau$. However, we see the opposite effect on compliance: While the results are relatively weak, they are bimodal, with a substantially greater probability density at higher compliance treatment effects for white than for nonwhite beneficiaries. One reason for this could be the language barrier for non-English-speaking applicants.
Figure 2: Comparison of average group outcomes under budget allocations for targeted treatment of different beneficiary shares ($E[Y(\pi_{\%})\mid A=a]$) with self-selection ($E[Y(\pi_{0})\mid A=a]$) or no-reminder ($E[Y(0)\mid A=a]$) status quo.
@@
The last plot shows the heterogeneity in the encouragement effect: When we combine the treatment and compliance effects, we see that the disparities in the compliance effects result in the distribution of the heterogeneous encouragement effect for white beneficiaries being wider than that for nonwhite ones. On the basis of such results, resource-constrained optimal rules would tend to target outreach to white individuals because of their higher “nudgeability". This could be concerning because, in some sense, the efficacy of recommendations can be changed by means of tactical redesign of the recommendation: more intensive outreach, pairing with a benefits counselor, etc. Koenecke et al. (2023) study efficient vs. equitable budget allocation for SNAP outreach via public surveys and find broad public support for equitable allocations, with potentially higher spending for targeting of harder-to-reach minority groups, for example.
Our final analysis considers equity in allocation efficiency, specifically the lift or improvements to decisions around who receives recommendations. In Figure˜3, we evaluate the lift in different performance metrics conditional on protected group and conditional on the group’s being encouraged under a resource-constrained policy—the y-axes show $\mathbb{E}[c(Y(\pi)))\mid A=a,\pi^{*}(X)=1],\mathbb{E}[(Y(1)-Y(0))\mid A=a,\pi^{*}(X)=1],\mathbb{E}[(T(1)-T(0))\mid A=a,\pi^{*}(X)=1]$. The x-axis ranges over thresholds on the conditional encouragement effect, $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\times\tau(X,A)$. Thus, the x-axis explores the range of the population covered under the budget allocation, with the budget averaged over the full population on the left-hand side and encouragement restricted to few individuals on the right-hand side. This represents the range of resource-constrained rules under a naive ITT approach. We do find, for large ranges of moderate budgets, that the naive efficient allocation results in disparities in average outcomes for white and nonwhite beneficiaries. The next two figures on the right disaggregate the individual encouragement effect into the heterogeneous causal effect $\tau$ of interview attendance and the compliance effect $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))$ of text message reminders on interview attendance. Consistent with the earlier results on the strong positive rank correlation of $\tau$ and $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))$, both plots are nearly nondecreasing as we threshold based on increasing individual encouragement scores $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\times\tau$. However, as we also saw earlier, this correlation is weaker for nonwhite beneficiaries, and we can observe a small region of nonmonotonic behavior.
@@
Comparing these two plots, we see that the disparities in naive ITT-based allocation are in fact driven almost entirely by disparities in access, which would persist under a large range of budgets. While leveraging disparities in Spearman’s correlation to diagnose disparities in resulting policy values is scale-free, Figure˜3 exactly highlights the magnitude of the conditional covariance gap.
Disaggregating our analysis by looking separately at treatment vs. compliance effects and at covariate-heterogeneous effects allows us to quantify access gaps that would persist even under efficient allocations. Isolating these sources supports firms in what to do next to reduce disparities, even if it may not be the fairness-constrained methods we next study in generality. Considering treatment effects and compliance effects separately is crucial since they imply focusing on different strategies moving forward, or because one may be outside of the firm’s control. For example, in the SNAP case study, only outreach and administrative (like how to schedule interviews) mechanisms have room for freedom for local organizations.
5 Method
First, we establish causal identification of the estimands via regression adjustment. Causal identification rewrites the causal estimand in terms of probability distributions estimable from data. The argument follows by our applying the conditional exclusion restriction and consistency but, crucially, does not rely on overlap. We also first consider a special type of fairness constraint, resource parity, and characterize optimal decisions.
Proposition 2 (Regression adjustment identification).
$\mathbb{E}[c(\pi,T(\pi),Y(\pi))]$ $\textstyle=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mu_{t}(X)p_{t\mid r}(X)]$
Resource parity–constrained optimal decision rules
@@ 5 Method
$V_{\epsilon}^{*}=\min_{\pi}\;\{\mathbb{E}[c(\pi,T(\pi),Y(\pi))]\colon\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]\leq\epsilon\}$ (1)
Enforcing absolute values, etc., follows in the standard way. Not all values of $\epsilon$ may be feasible; in the appendix, we give an auxiliary program for computing feasible ranges of $\epsilon$. We first characterize a threshold solution when the policy class is unconstrained.
Proposition 3 (Threshold solutions under resource constraints).
Define $L(\lambda,X,A)=$
$(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\left\{\tau(X,A)+\frac{\lambda}{p(A)}(\mathbb{I}\left[A=a\right]-\mathbb{I}\left[A=b\right])\right\}+\lambda(p_{1\mid 0}(X,a)-p_{1\mid 0}(X,b))$
. Then, $\lambda^{*}\in\arg\min_{\lambda}\;\mathbb{E}[L(\lambda,X,A)_{+}]$ and $\pi^{*}(x,u)=\mathbb{I}\{L(\lambda^{*},X,u)>0\}.$ If instead $d(x)$ is a function of covariates $x$ only, $\lambda^{*}\in\arg\min_{\lambda}\;\mathbb{E}[\mathbb{E}[L(\lambda,X,A)\mid X]_{+}]$ and $\pi^{*}(x)={\mathbb{I}\{\mathbb{E}[L(\lambda^{*},X,A)\mid X]>0\}}.$
Establishing this threshold structure (which follows by duality of infinite-dimensional linear programming) allows us to provide a generalization bound argument.
Proposition 4 (Policy value generalization).
Assume that the nuisance models $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0}]^{\top},\eta\in\mathcal{F}_{\eta}$ are consistent and well specified with finite Vapnik–Chervonenkis (VC) dimension $v_{\eta}$ over the product function class $\mathcal{F}_{\eta}$. Let $\Pi=\{\mathbb{I}\{\mathbb{E}[L(\lambda,X,A;\eta)\mid X]>0\colon\lambda\in\mathbb{R};\eta\in\mathcal{F}_{\eta}\}.$
$\sup_{\pi\in\Pi,\lambda\in\mathbb{R}}\left|(\mathbb{E}_{n}[\pi L(\lambda,X,A)]-\mathbb{E}[\pi L(\lambda,X,A)])\right|=O_{p}(n^{-\frac{1}{2}})$
@@
Proposition 5 (Variance-reduced estimation).
$\textstyle V(\pi)$ $=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\frac{\mathbb{I}\left[R=r\right]}{e_{r}(X)}(\mathbb{I}[T=t]c_{rt}(Y)-\mu_{t}(X)p_{t\mid r}(X))+\mu_{t}(X)p_{t\mid r}(X)\right\}\right]$
$\mathbb{E}[T(\pi)]$ $=\textstyle\sum_{r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\frac{\mathbb{I}\left[R=r\right]}{e_{r}(X)}(T(r)-p_{1\mid r}(x))+p_{1\mid r}(x)\right\}\right]$
We retain the full expression rather than simplifying (as appears in Qiu et al., 2021) since the doubly robust estimation of constraints changes the Lagrangian. For example, for regression adjustment, it is clearer in Proposition˜10 how constraints affect the optimal decision rule.
5.1 Robust estimation with treatment overlap but without recommendation overlap
When the recommendations are, e.g., the high risk/low risk labels from binary classifiers, the overlap assumption may not be satisfied since the algorithmic recommendations are deterministic functions of covariates. However, note that identification in Proposition˜2 requires only SUTVA, consistency, and the exclusion restriction.
A naive approach based on parametric extrapolation is to estimate $p_{1\mid 1}(X)$, treatment responsivity, on the observed data and simply use the parametric form to extrapolate to the full dataset. (In Appendix˜B, we describe the variance reduction that can be possible). On the other hand, parametric extrapolation is generally unsatisfactory because the conclusions will be driven by the model specification rather than by observed data. Nonetheless, it can provide a starting point for robust extrapolation of structurally plausible treatment response probabilities.

### v5

@@ 4.2 Case study: Text message reminders for SNAP recertification
We illustrate the above diagnostic analysis of compliance and treatment benefit in a case study drawing on a randomized-controlled trial of text message reminders for SNAP recertification. We have two purposes: first, we ask whether racial disparities in the encouragement score $s(X)=\tau(X)\kappa(X)$ are driven primarily by treatment benefit or by compliance. Second, we illustrate how these differences translate into disparities under the actual threshold policies induced by naive ITT targeting.
@@
Each year, households receiving SNAP (food assistance) must complete a short recertification interview to verify continued eligibility. Missing this interview results in termination of benefits, even for households that otherwise remain eligible. We study a pilot program in San Francisco in which clients could opt in to receive a text-message reminder about the recertification deadline (Homonoff and Somerville 2021). The reminder encouraged participants to complete the required interview—our effective treatment—so that their benefits would continue. Since many who fail to recertify are in fact eligible, increasing interview attendance often improves benefit continuity and household welfare.
@@
We reinterpret the text message reminder as an encouragement $R$ affecting the take-up of treatment $T$ (interview completion), which in turn influences the outcome $Y$ (next-year SNAP benefits). We focus on heterogeneity in both the reminder’s compliance effect and the treatment effect of interview completion, with particular attention to racial disparities in their alignment.
Step 1: Descriptive diagnostics: heterogeneity in treatment benefit and compliance, and their alignment.
We begin by comparing the groupwise distributions within each race (binarized to white $A=0$ or nonwhite $A=1$) in Figure 1 of the heterogeneous treatment effect $\tau(X)$ the heterogeneous compliance effect $\kappa(X)$ and the encouragement score which is their product and which gives the ranking for optimal budget-constrained encouragement targeting rules.
Figure 1 shows these descriptive diagnostics. The first plot shows the treatment effect heterogeneity: The distribution is wider for nonwhite than for white beneficiaries; i.e., it has more mass at higher magnitudes of $\tau$ and less at lower magnitudes of $\tau$. However, we see the opposite effect on compliance: While the results are relatively weak, they are bimodal, with a substantially greater probability density at higher compliance treatment effects for white than for nonwhite beneficiaries. One reason for this could be the language barrier for non-English-speaking applicants. The last plot shows the heterogeneity in the encouragement effect: the distribution of the heterogeneous encouragement effect for white beneficiaries is wider than that for nonwhite ones. These diagnostics show that group-level differences in nudgeability dominate in disparities in the encouragement score.
Another source of potential disparities in improvement from personalized targeting arises from differential alignment between treatment take-up and efficacy. We quantify this alignment with the Spearman’s $\rho_{SP}$ rank correlation coefficient. Among white beneficiaries only, the rank correlation coefficient between heterogeneous treatment and compliance effects is $0.68$, while it is $0.45$ for nonwhite beneficiaries. While both correlations are statistically significantly positive, the relationship is weaker for nonwhite than for white beneficiaries. This means that non-whites who benefit the most from treatment may not be the most responsive to nudges into it. Spearman correlation is not itself the conditional covariance term in the improvement decomposition—it is rank-based, scale-free, and not conditioned on selection—but it is a useful threshold-free diagnostic proxy for disparities in policy value and/or improvement due to differential alignment.
@@
Step 2: Investigate disparities under naive ITT targeting.
We next investigate how these differences result in disparities in the actual allocations induced by naive ITT targeting. To do so, we consider threshold policies as in Equation 2 of the form $\pi_{\%}(X)=\mathbb{I}\{s(X)\geq\%\text{percentile of }s\},$ since they solve naive-ITT budgeted targeting.
Whereas Step 1 uses unconditional, threshold-free diagnostics to assess whether treatment benefit and compliance are differently aligned across groups, Step 2 evaluates the conditional means of these same quantities within the sets actually selected by naive ITT threshold rules, which are exactly (proportional to) the objects governing the groupwise improvement decomposition. Figure 2 connects the previous mechanism-based diagnostic to investigate who is selected by naive ITT targeting under different policies.
For each threshold, we look within each group at those who would be selected and plot the average encouragement score, treatment benefit, and compliance (y-axis, respectively). We sweep over score thresholds on the x-axis to sweep over all possible budgets.
The left panel shows the average encouragement score among selected individuals by group. For large ranges of moderate budgets, the selected white applicants have larger encouragement score than non-white applicants. The middle plot illustrates the treatment benefit $\tau$, which is broadly similar within both group’s selected sets, so disparities in average outcome are not driven by differences in how much individuals stand to gain from interview completion. By contrast, the right panel shows a clearer gap, persisting across budgets, in the average compliance effect $\mathbb{E}[(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\mid\pi^{*}=1,A]$.
@@
Comparing these panels, we see that disparities in improvement naive ITT-based allocation are in fact driven almost entirely by disparities in access, which would persist under a large range of budgets.
Step 3: Interpretation of potential disparities, if present.
A central concern is that worse-off groups may benefit more from treatment yet be less responsive to low-touch encouragement because of administrative and resource barriers. In SNAP, the former is plausible given diminishing marginal utility, while the latter is consistent with the burdens of recertification, scheduling, childcare, and transportation. This distinction matters because responsiveness is potentially actionable: organizations can redesign outreach intensity, navigation support, or interview logistics. Koenecke et al. (2023) likewise find public support for more equitable SNAP outreach, including greater spending to reach harder-to-reach groups.
By separating treatment efficacy from responsiveness, our audit identifies access gaps that would persist even under outcome-efficient allocations. next step. Knowing whether limits to improvements from personalized targeting stem from disparities in compliance or treatment efficacy clarifies relevant next steps. In SNAP, that next step lies in outreach and administrative redesign rather than the underlying benefit itself. However local administrators and nonprofits have far more room to design supplemental outreach processes affecting compliance, than they do to modify the centralized policy eligibility rules that underlie treatment efficacy.
Step 4: From audit to redesign: quantifying the gains and limits of personalized targeting in SNAP.
We use off-policy evaluation and optimization to quantify the improvements and limits of personalized targeting in SNAP, illustrating our how prior diagnosed disparities result in potential equity concerns in group-differential limits to improvement from targeting, due to persistent access barriers.
Figure 3: Comparison of average group outcomes under budgeted allocations (solid lines) for targeted treatment of different beneficiary shares ($E[Y(\pi_{\%})\mid A=a]$) with self-selection ($E[Y(\pi_{0})\mid A=a]$) or no-reminder ($E[Y(0)\mid A=a]$) status quo.
We quantify the extent of potential improvements from automatic enrollment in reminders (recommend to all) or, under potential budget constraints, budget-optimal targeted allocations. In Figure 3, the y-axis measures different groupwise or average downstream values: red for the group $A=1$ (nonwhite), $\mathbb{E}[Y(\pi)\mid A=1]$; purple for the entire population, $\mathbb{E}[Y(\pi)]$; and blue for $A=0$ (white), $\mathbb{E}[Y(\pi)\mid A=0]$. Next, within each set of colored lines, we compare the following: the solid line for budget-optimal policies $\mathbb{E}[Y(\pi_{\%})]$, the densely dotted line treating all ($E[Y(1)\mid A=a]$), the dashed line indicating the observed self-selection, and the sparsely dotted line treating no one ($E[Y(0)\mid A=a]$). The x-axis ranges from treating everyone to no-one by ranging over encouragement score percentile thresholds, hence sweeping over all possible budget restrictions. This figure also connects back to the earlier fairness decomposition. For each group $a$, the groupwise improvement $\mathbb{E}[Y(\pi_{b})\mid A=a]-\mathbb{E}[Y(0)\mid A=a]$ can be read off the plot as the difference between solid and sparsely-dotted lines. Observe that the magnitude of these improvements differs across groups, exactly due to our previously observed disparities in compliance and alignment between treatment take-up and efficacy.
Overall, additional reminders weakly improve outcomes relative to no reminders, and relative to currently observed self-selection. The main exception is the white subgroup, for whom self-selection modestly improves outcomes. By contrast, for budgets covering at least roughly $25\%$ of the population, targeted allocation substantially improves group-average outcomes relative to both current self-selection and the no-reminder baseline. Thus, in terms of group-level outcomes, targeted reminder allocation improves upon the status quo. Nonetheless, limits persist as to the extent of gains from low-touch outreach alone, suggesting that additional investment in outreach resources beyond text reminders to ensure that those who may benefit the most from program enrollment can overcome access barriers or burdens to do so. Additional research would be needed to ascertain root causes and whether the observed disparities in compliance reflect behavioral inattention, trust in government, or structural barriers to compliance with administrative burdens matters for policy implications.
5 Policy Optimization under Access Constraints
Here we move from diagnosis to decision rules. We first characterize the structure of optimal fairness-aware encouragement policies in the population—showing they remain threshold rules on appropriate scores or Lagrangian scores—and then use those moment representations to motivate the estimation and constrained policy-learning procedures that follow.
Although our motivating decision problem is multi-objective, the formal statistical analysis below focuses on the primary outcome criterion $V(\pi)=\mathbb{E}[Y(\pi)]$. Fairness, access, and budget considerations enter through separately estimated constraint moments, and application-specific scalarizations can be formed from those same moments when desired.
Proposition 5.1 (Group-specific thresholds under demographic parity. )
Suppose the distribution of $s(x,a)$ is continuous (no atoms). Consider optimizing encouragement effects under a $\beta$-fraction budget and groupwise budget parity:
$\pi_{DP}^{*}(x,a)\in\max_{\pi(x,a)}\{\mathbb{E}[Y(\pi)]\colon\mathbb{E}[\pi(X,A)]\leq\beta,\left|\mathbb{E}[\pi(X,A)\mid A=a]-\mathbb{E}[\pi(X,A)\mid A=b]\right|=0\}$ (3)
Define the $\beta$-quantile for group $a$’s score: $t_{\beta}^{a}\coloneqq\max\!\left\{\inf\{z:P(s(X,a)\leq z)\geq 1-\beta\},\,0\right\}$. Then the optimal policy comprises group-dependent thresholds at the top-$\beta$ quantile of the encouragement score within each group (so long as it is positive):
$\pi_{DP}^{*}(x,a)\;=\;\mathbb{I}[s(x,a)\geq t_{\beta}^{a}].$
That is, ensuring parity in encouragement budget by group is achieved by ranking within the score distributions in each group.
Treatment parity–constrained optimal decision rules
@@ 5 Policy Optimization under Access Constraints
$\max_{\pi}\;\{\mathbb{E}[Y(\pi)]\colon\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]\leq\epsilon\}$ (4)
We discuss the one-sided constraint in the analysis for simpler discussion; enforcing absolute values, etc., follows in the standard way. Not all values of $\epsilon$ may be feasible, and the amount of room for reducing disparities in treatment depends on the heterogeneity in treatment response. The feasible range of treatment-parity disparities is an interval $\inf_{\pi}\left|\Delta_{T}\right|,\sup_{\pi}\left|\Delta_{T}\right|]$ that can be pre-calculated, so that a given treatment disparity constraint $\Delta_{T}\leq\epsilon$ is feasible iff $\epsilon$ lies within the interval.
We first characterize a threshold solution when the policy class is unconstrained.
Proposition 5.2 (Threshold solutions under treatment-parity constraints)
Let $\kappa(X,A):=p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A)$, $c_{0}:=\mathbb{E}[p_{1\mid 0}(X,A)(\tfrac{\mathbb{I}\left[A=a\right]-\mathbb{I}\left[A=b\right]}{p(A)})]$, and
$\ell(\lambda,X,A):=\kappa(X,A)\bigl\{\tau(X,A)-\lambda\tfrac{\mathbb{I}\left[A=a\right]-\mathbb{I}\left[A=b\right]}{p(A)}\bigr\}.$
Suppose (4) is strictly feasible. Then there exists $\lambda^{*}\in\arg\min_{\lambda\geq 0}\bigl\{\mathbb{E}[\ell(\lambda,X,A)_{+}]+\lambda(\epsilon-c_{0})\bigr\}$ such that $\pi^{*}(x,u)=\mathbb{I}\{\ell(\lambda^{*},x,u)>0\}$. If instead $\pi$ is restricted to covariates only, $\lambda^{*}\in\arg\min_{\lambda\geq 0}\bigl\{\mathbb{E}[\mathbb{E}[\ell(\lambda,X,A)\mid X]_{+}]+\lambda(\epsilon-c_{0})\bigr\}$ and $\pi^{*}(x)=\mathbb{I}\{\mathbb{E}[\ell(\lambda^{*},X,A)\mid X=x]>0\}$.
Establishing this threshold structure (which follows by duality of infinite-dimensional linear programming) allows us to provide a generalization bound argument. We make a standard assumption that the functions we use for estimation are learnable, e.g. have finite VC dimension.
{assumption}
[Learnable nuisance functions] Assume that the nuisance models $\eta=[p_{1\mid 0},p_{1\mid 1},\mu_{1},\mu_{0}]^{\top},\eta\in\mathcal{F}_{\eta}$ are consistent and well specified with finite Vapnik–Chervonenkis (VC) dimension $v_{\eta}$ over the product function class $\mathcal{F}_{\eta}$.
Proposition 5.3 (Policy value generalization)
Let $\Pi=\{\mathbb{I}\{\mathbb{E}[\ell(\lambda,X,A;\eta)\mid X]>0\colon\lambda\in\mathbb{R};\eta\in\mathcal{F}_{\eta}\}.$
$\sup_{\pi\in\Pi,\lambda\in\mathbb{R}}\left|(\mathbb{E}_{n}[\pi\ell(\lambda,X,A)]-\mathbb{E}[\pi\ell(\lambda,X,A)])\right|=O_{p}(n^{-\frac{1}{2}})$
@@
Proposition 5.4 (Variance-reduced estimation)
$\textstyle V(\pi)$ $=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\frac{\mathbb{I}\left[R=r\right]}{e_{r}(X)}(\mathbb{I}[T=t]Y-\mu_{t}(X)p_{t\mid r}(X))+\mu_{t}(X)p_{t\mid r}(X)\right\}\right]$
$\mathbb{E}[T(\pi)]$ $=\textstyle\sum_{r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\frac{\mathbb{I}\left[R=r\right]}{e_{r}(X)}(T-p_{1\mid r}(x))+p_{1\mid r}(x)\right\}\right]$
We retain the full expression rather than simplifying (as appears in Qiu et al. 2021) since the doubly robust estimation of constraints changes the Lagrangian. For example, for regression adjustment, it is clearer in Proposition 5.2 how constraints affect the optimal decision rule.
5.1 Extension: Algorithmic advice with treatment, but not recommendation overlap
When the recommendations are, e.g., the high risk/low risk labels from binary classifiers, the overlap assumption may not be satisfied since the algorithmic recommendations are deterministic functions of covariates. However, note that identification in Proposition 3.1 requires only SUTVA, consistency, and the exclusion restriction.
A naive approach based on parametric extrapolation is to estimate $p_{1\mid 1}(X)$, treatment responsivity, on the observed data and simply use the parametric form to extrapolate to the full dataset. Though this is generally unsatisfactory by itself, it is the starting point for our robust extrapolation methods.